OpenStudy (anonymous):

Find the volume of a cylinder with a diameter of 8 inches and a height that is three times the radius. Use 3.14 for pi and round your answer to the nearest hundredth. (Hint: You may only enter numerals, decimal points, and negative signs in the answer blank

OpenStudy (anonymous):

help @UnkleRhaukus

OpenStudy (tkhunny):

Diameter = 8" Radius = ???

OpenStudy (anonymous):

its says 3 times the radius

OpenStudy (tkhunny):

Answer the question, please? Given that the Diameter is 8", what is the Radius? There is a relationship between Diameter and Radius.

OpenStudy (anonymous):

4 i think

OpenStudy (tkhunny):

Since Diameter = 2* Radius, it is 4". No more just "thinking". Substitute with "knowing". Okay, Diameter = 8" Radius = 8"/2 = 4" Height = ??

OpenStudy (anonymous):

would it be 12

OpenStudy (anonymous):

or 64

OpenStudy (tkhunny):

I would like for you to abandon your question marks. Be SURE! Diameter = 8" Radius = 8"/2 = 4" Height = 3*4" = 12" Reread the problem statement until you are SURE. No guessing. Volume = ??

OpenStudy (anonymous):

603.2

OpenStudy (tkhunny):

It helps if you write stuff out. This allow you to check for errors and to explain and understand your work. Diameter = 8" Radius = 8"/2 = 4" Height = 3*4" = 12" Volume = \(\pi r^{2} h = 3.14\cdot(4")^{2}\cdot 12" = 3.14\cdot 192 = 602.88\) Everything I did is right there on the screen - the WHOLE presentation! I recommend this. Because you have not been quite that generous with your calculations, I cannot tell why you developed a rounding problem. I suspect you did not use the specified value for \(\pi\). You used the \(\pi\) button on your calculator, didn't you? If this is the case, you should have gotten 603.18578948924, which, when rounding to the hundreths place is 603.19. Recap: 1) Be more clear and open. Write things down. 2) If the problem statement says use 3.14 for \(\pi\), do NOT use the \(\pi\) button on your calculator. 3) Your rounding is not quite as it should be. 4) Reread the problem statement one last time before you think you are done. Make SURE you have fulfilled all the requirements and complied with whatever prescription is given.

OpenStudy (anonymous):

Thank you